# HF radar estimation of ocean wave parameters: second-order Doppler spectrum versus Bragg wave modulation approach

### v1

@@ 3.2 First-order Bragg peak method
where $t$ is time; $R^{\pm}$ and $a^{\pm}$ are the complex amplitudes with random phase of the long and Bragg waves respectively, $\omega_{p}$ is the peak frequency of the long wave and $\omega_{u}=4\pi u/\lambda$ is the additional frequency shift due to the radial surface current $u$. We expect $H_{s}$ to be proportional to the variance of the long wave modulation, i.e:
$H_{s}\sim\mathrm{Var}{(R^{+})}+\mathrm{Var}{(R^{-})}$ (6)
@@
Note that the oscillation factor due to long wave modulation “survives” the radar cell averaging process because the wavelength of the longest waves is not negligible compared to the radar range resolution. For typical scales of $75$ m for the former and $1.5$ km for the latter, about $20$ long waves pass through a given radar cell, so that the mean surface level within that patch oscillates in phase with the long wave with a damping factor of about $1/20$. The amplitude of this damped oscillation remains large compared to the amplitude of the Bragg wavelength (because the ratio of the peak wavelength to the Bragg wavelength is $75/9.3\sim 8$ and the ratio of the corresponding amplitude is $\sim 8^{2}$ for a $k^{-4}$ wave spectrum). In addition, there are also wave groups of comparable or larger size to the radar cell. These groups are responsible for very low frequency modulation of the radar signal.
Since both the peak wave frequency $\omega_{p}$ and the current-induced frequency shift $\omega_{u}$ are much smaller than the Bragg frequency $\omega_{B}$, we can rewrite eq. (5) as the sum of an analytical signal $s_{a}$ (respectively, anti-analytical signal $s_{aa}$) that contains only the positive frequency side (respectively, negative frequency side) of the Bragg peaks:
$s(t)=s_{a}(t)+s_{aa}(t)=A^{+}(t)e^{i\omega_{B}t}+A^{-}(t)e^{-i\omega_{B}t},$ (7)
@@
$A^{+}(t)=(R^{+}a^{+}e^{i\omega_{p}t}+R^{-}a^{+}e^{-i\omega_{p}t})e^{i\omega_{u}t}$ (8)
and
$A^{-}(t)=(R^{+}a^{-}e^{i\omega_{p}t}+R^{-}a^{-}e^{-i\omega_{p}t})e^{i\omega_{u}t}$ (9)
Assuming the complex amplitudes $R^{\pm}$ and $a^{\pm}$ to be uncorrelated due to random phase shifts between approaching and receding waves, we obtain that the variance of the analytic signal combined with the variance of the anti-analytic signal correlates with the variance of the ocean’s free surface associated with long waves we have:
$\mathrm{\langle\lvert A^{+}\rvert^{2}\rangle}+\mathrm{\langle\lvert A^{-}\rvert^{2}\rangle}\sim\mathrm{\langle\lvert R^{+}\rvert^{2}\rangle}+\mathrm{\langle\lvert R^{-}\rvert^{2}\rangle},$ (10)
@@
Now from eqs (8) and (9) it appears that the absolute amplitudes $\lvert A^{\pm}\rvert$ can be extracted from the complex radar signal by taking the absolute value of its analytical and anti-analytical components:
$\lvert A^{+}(t)\rvert=\lvert s_{a}(t)\rvert,\ \lvert A^{-}(t)\rvert=\lvert s_{aa}(t)\rvert$ (11)
By construction the variables $\lvert s_{a}\rvert$, $\lvert s_{aa}\rvert$ and $\lvert R^{\pm}\rvert$ can be assumed to be Rayleigh-distributed random variables (as the modulus of complex Gaussian random variables). For this type of variable, the variance is a constant proportion of the second moment (i.e. $\mathrm{Var}{(X)}=\alpha\mathrm{\langle\lvert X\rvert^{2}\rangle}$ for some $\alpha>0$, here $\alpha=(1-\pi/4)$). It follows from (eq. 6) that the $H_{s}$ can be obtained as:
$H_{s}^{2}=C(\langle\mathrm{Var}{(\lvert s_{a}\rvert)}\rangle+\langle\mathrm{Var}{(\lvert s_{aa}\rvert)}\rangle)$ (12)
@@
where $C$ is some calibration factor that must be set for each range and azimuth. This factor is calculated as the proportional coefficient between the mean value for the whole $2020$ year of the buoy $H_{s}$ and the mean value of the sum of $\lvert A^{+}\rvert$ and $\lvert A^{-}\rvert$ variances for the same period.
Figure 5: a) Real part of the received radar signal time series for the range $10$ and azimuth $75$ after beamforming (black) with the modulus of its analytical and anti-analytical signal (red) for September $27$, $2020$, at $23$h where $H_{s}$ and $T_{p}$ from CANDHIS are $2.93\ m$ m and $7.5\ s$, respectively. b) Modulus of the analytical signal for the same time. c) Power spectrum from the sum of the analytical and anti-analytical signals ($\sigma_{\lvert sa\rvert}(\omega)+\sigma_{\lvert saa\rvert}(\omega)$) for the same time. The common wave frequencies in the Mediterranean Sea are marked with a red rectangle where the maximum power spectrum is related to the peak frequency ($\omega_{p}$)
Since the amplitude modulation $\lvert A^{\pm}(t)\rvert$ follows the variation of long waves, the peak frequency ($\omega_{p}$) can be identified as the dominant frequency in the associated power spectrum:
$\omega_{p}=\mathrm{Argmax}(\sigma_{\lvert sa\rvert}(\omega)+\sigma_{\lvert saa\rvert}(\omega)),$ (13)
@@
where $\sigma_{\lvert sa\rvert}$ and $\sigma_{\lvert saa\rvert}$ are the power spectra associated with the absolute value of the analytical and anti-analytical radar signals respectively. The $T_{p}$ can be derived from the relationship $T_{p}=1/\omega_{p}$. Figure 5 a) shows an example of the complex radar signal time series recorded on the antenna receiving array after beamforming processing. A slow amplitude modulation is visible, marked by the red lines. Part of this modulation is due to the beat frequency caused by the shift of the surface current from the Bragg frequency (see Guérin and Grilli, (2018)) and another part is due to the long wave groups (eqs. (8) and (9) ). This last term can be isolated by taking the modulus of the analytical/anti-analytical signal (see figure 5, b)), which has the effect of canceling out the Bragg component and the current-induced frequency shift. The peak wave frequency is sought at the maximum of the corresponding power spectrum (Figure 5, c)) in the range $0.1$ to $0.4$ Hz, which contains the dominant wave frequencies for the Mediterranean Sea. Only the positive part of the power spectrum is shown as it is symmetric.
@@ 4 Results and Discussion
This section is divided into the estimation of $H_{s}$ on the one hand and $T_{p}$ on the other since the precision of the estimation of these two variables is quite different. In this way, the problems and limitations encountered in the estimation of each variable can be discussed separately.
4.1 Significant wave height Estimation
The results for $H_{s}$ for the whole year $2020$, obtained using data from the Fort Peyras station only, are shown in Figure 6. Only minimal post-processing has been done by removing the obvious outliers $H_{s}>4$ m and $T_{p}>10$ s in the hourly estimates and performing a moving median average over 3 consecutive range bins (i.e. 4.5 km) and 5 consecutive azimuths (i.e. 5 degrees) to smooth the data and fill in some gaps. These results are shown along the azimuth $40$, as defined in Figure 1, at different distances from the radar station: in the short range (radar cell $8$ at $12$ km, hereafter referred to as $r8$); in the medium range (radar cell $20$ at $30$ km, $r20$); in the long-range (radar cell $30$ at $52.5$ km, $r30$). As can be seen, the results obtained using the SODS method are in good agreement with the data collected by the CANDHIS buoy. However, the accuracy of this method decreases with increasing distance from the radar. When we reach a distance of $20$ ($30$ km from the radar), it is no longer possible to estimate $H_{s}$ because the second-order level of the energy spectrum becomes indistinguishable from the noise. For this reason, Figure 6, c (in the range $35$ to $52.5$ km from the radar) does not show the results obtained with the SODS method. On the contrary, the estimate based on the FOBP is still relevant, if not accurate, at a greater distance.
Figure 6: Time series of $H_{s}$ (m) along the azimuth $40$ at a) range $8$ ($12$km), b) range $20$ ($30$km) and c) range $35$ ($52.5$km). The black line represents the data from the CANDHIS buoy located at Porquerolles, the grey line is the results of the WaveWatch III model run in the area, the red and blue lines correspond to the $H_{s}$ obtained using the SODS and the FOBP method from the radar data, respectively.
Table 1: Statistical performances of the SODS and FOBP methods measured over the year $2020$ along the central azimuth $a40$ at cell range $8$, $20$, and $35$ ($12$, $30$, and $52.5$ km, respectively) and compared with the ground truth of the CANDHIS buoy. The first index is the median-product correlation coefficient ($R^{\ast}$), the second index is the RMS difference (RMSE) and the last index is the difference of variances ($\Delta V$) in the time series.
$R^{\ast}$ RMSE ($m$ or $s$) $\Delta V$ ($m^{2}$ or $s^{2}$)
SODS FOBP SODS FOBP SODS FOBP
$H_{s}$ r8 $0.92$ $0.77$ $0.39$ $0.66$ $0.002$ $0.15$
r20 $0.88$ $0.72$ $0.66$ $0.61$ $0.33$ $0.06$
r35 $0.59$ $0.64$ $0.03$
$T_{p}$ r8 $0.56$ $0.69$ $1.60$ $1.67$ $1.71$ $0.11$
r20 $0.65$ $0.66$ $1.39$ $1.57$ $1.43$ $0.43$
r35 $0.46$ $1.73$ $0.92$
In order to compare both methods more extensively, the $H_{s}$ scatterplots for the $r8$ and $r20$ ranges (those for which SODS and FOBP results are available) along the $40$ azimuth are shown in Figure 7. It can be seen how for the SODS method the estimates are less scattered, although it overestimates the value of $H_{s}$ in all its value ranges, being this more evident farther away from the radar ($r20$). On the other hand, although the FOBP method shows a larger spreading, the observed deviation from the reference values is more symmetrical, neither overestimating nor underestimating the buoy $H_{s}$ value. It is also worth noting that the $H_{s}$ estimation with the FOBP method is more stable as we move away from the radar; we can see how for $r8$ and $r20$ similar shapes are obtained.
@@ 4.1 Significant wave height Estimation
Figure 7: Scatter plot of radar $H_{s}$ from (a and c) SODS method and (b and d) FOBP method versus $H_{s}$ from CANDHIS buoy where each point is colored by the spatial density of nearby points. The spatial locations are along the $40$ azimuth and the ranges $8$ (a and b) first row, and $20$ (c and d) second row.
The respective performances of the $2$ methods have been quantified with a few statistical indices, using the CANDHIS buoy data as ground truth. The first index is the median product correlation coefficient $R^{\ast}$ (Shevlyakov, (1997)), which is more robust than Pearson’s correlation coefficient in series with noise and jumps. This index has already been used in the context of HF radars to assess surface currents (Dussol et al., (2022)) and we found it useful in the context of wave estimation. It is defined as
$R^{\ast}=\frac{\mathrm{Med}^{2}(\lvert a\rvert)-\mathrm{Med}^{2}(\lvert b\rvert)}{\mathrm{Med}^{2}(\lvert a\rvert)+\mathrm{Med}^{2}(\lvert b\rvert)}$ (14)
where:
$\begin{split}a&=H_{s\;radar}-\mathrm{Med}(H_{s\;radar})+\left(H_{s\;CANDHIS}-\mathrm{Med}(H_{s\;CANDHIS})\right);\\ b&=H_{s\;radar}-\mathrm{Med}(H_{s\;radar})-\left(H_{s\;CANDHIS}-\mathrm{Med}(H_{s\;CANDHIS})\right);\end{split}$ (15)
The second index is the Root Mean Square Error (RMSE) between the directional $H_{s}$ data from the radar and those from the buoy, which are taken as reference values.
Note that the RMSE is not a perfect indicator of accuracy as it can be artificially increased by the presence of outliers or, conversely, kept artificially low by taking a constant median of the time series. It is therefore useful to complement this quantity with an additional statistical index that compares the variability of the respective time series. This can be done simply by evaluating the difference $\Delta V$ of their respective variances. The results for the $2$ methods and the $3$ statistical indices are summarised in table 1 for the locations shown in figure 6 ($a40$ and $r8$, $r20$ and $r35$). In the short and medium range, the SODS estimate remains highly correlated with the buoy measurement ($R^{\ast}\simeq 0.9$), but its RMSE deteriorates from $0.39$ m to $0.66$ m, as does the difference in variances ($\Delta V$), which increases by two orders of magnitude; this difference is visually illustrated in Figure 6, b). On the contrary, the performance of the FOBP, although not optimal, remains remarkably stable with increasing range ($R^{\ast}\simeq 0.6-0.8$ and RMSE0.65similar-to-or-equalsabsent0.65$\simeq$ 0.65≃ 0.65 m). In this case, the $\Delta V$ remains more constant and even decreases as we move away from the radar. Note that the obtained RMSE values should be compared with the mean value of $H_{s}$ observed in the Toulon region (Figure 1a), which ranges from about 0.9 m to 1.2 m with increasing distance, and with the intrinsic RMS difference between the $H_{s}$ at the radar and buoy location, which ranges from $0.1$ m to $0.4$ m (Figure 1b).
A spatial distribution of these statistics in the Toulon radar area is shown in figure 8. It can be seen that the above statements are true not only along the azimuth $40$ but over the whole radar coverage. It should be noted that the SODS method is very sensitive to any electromagnetic interference that may be present in the area, and for this reason, the results obtained in the shadow of the Giens tombolo (east of the radar coverage) are of poorer quality. On the other hand, the FOBP method, based on the first-order Bragg signal, is less affected by such interference and remains satisfactory even in the vicinity of the Giens peninsula. However, the latter method suffers from another type of interference, namely the power grid artifact discussed in section 3.1, which distorts the received signal at distances of about $37.5$ and $45$ km from the radar site. These data have been excluded from the results.
Figure 8: Spatial maps of the $3$ statistical performance indicators for the radar measurement of $H_{s}$ obtained for the year $2020$ with the SODS (left column, a, c, e) and FOBP (second column, b, d, f) methods: 1) $R^{\ast}$ (a, b); 2) RMSE (c, d); 3) $\Delta V$ (e, f). The black arcs mark the ranges r8, r20, and r35 ($12$, $30$, and $52.5$ km) respectively, and the black radial indicates the azimuth used for the calculations in table 1. The SODS is limited to r35, while the FOBP extends well beyond. The white circular bands in the maps in the right column correspond to the few areas where the low-frequency component of the radar signal is distorted by a power grid artifact.
4.2 Peak wave period Estimation
Regarding the results obtained for $T_{p}$ with the classical SODS method, we can observe in Figure 9 and Table 1 worse performances, with a reduced correlation coefficient ($R^{\ast}\simeq 0.6$) and a relatively large RMS error (RMSE1.5similar-to-or-equalsabsent1.5$\simeq$ 1.5≃ 1.5 seconds, that is about $24\%$ of the mean peak period $6.14$ seconds). From Figure 9 it can be seen that the SODS radar estimate is not able to follow the full variability of $T_{p}$, especially towards the lowest values. This could be due to the reduced accuracy of the asymptotic formula (eq. 4) at low wind speeds, as suggested by recent numerical simulations (Guérin,, 2024). The $T_{p}$ estimated with the FOBP shows similar performances to SODS in terms of correlation coefficient and RMSE, even in the near range (Table 1). However, as can be seen from the plots in Figure 9, this new estimate has a much better ability to follow the fluctuations of $T_{p}$, even at its lowest values, and it is robust with increasing range. This is reflected in the difference in the variances of the time series, which is much smaller than with the SODS method (Table 1).
Figure 9: Time series of $T_{p}$ (s) along the azimuth $40$ at a) range $8$ ($12$km), b) range $20$ ($30$km) and c) range $35$ ($52.5$km). The black line represents the data from the CANDHIS buoy located at Porquerolles, the grey line is the results of the WaveWatch III model run in the area, the red and blue lines correspond to the $T_{p}$ obtained using the SODS and the FOBP method from the radar data.
The scatter plot analysis confirms these results (Figure 10). It is shown how the SODS method cannot capture $T_{p}$ less than $4$ s due to the limitation of distinguishing the second order of the spectrum from the noise, while the FOBP method, which does not have this limitation, can capture periods up to $2.5$ s. For $T_{p}$ the same happens as for the $H_{s}$ estimation, the SODS method, despite having a narrower shape, overestimates the buoy values, although in this case no deterioration is observed as we move away from the radar. As for the FOBP method, the shape obtained is symmetrical for high periods, while in the lower period range, this method underestimates the value of the buoy period. This may be due to the large number of wave periods of similar energy that can be found in a calm sea. In this situation, the FOBP method may give erroneous values, but close to the ground truth. This fact can be better explained in figure 5, c), where we can see that there are at least $3$ peaks at frequencies lower than $\omega_{p}$ with energy very close to that selected by the FOBP method. On the other hand, when high periods are found, we are usually in storm conditions where the waves are more extreme and more ordered in wavefronts, identified by a strong peak energy around $\omega_{p}$, making it easier to identify this value. For this reason, and although FOBP can solve for periods smaller than $4$s, this method is still at an early stage of development and may improve the estimation of $T_{p}$ at low period values in future work.
Figure 10: Same as Figure 7 for $T_{p}$.
The results highlighted with some radar cells ($r8$, $r20$, and $r35$) along the central azimuth ($a40$) have been confirmed with other radar cells (see figure 11). As for the $H_{s}$ estimation, the ranges are truncated at $30$ km for the SODS method, while for the FOBP method, the valid ranges extend up to $52.5$ km. The missing values around $19.5-24$km and $37.5-45$km (indicated by the white circle in figure 8.b, 8.d and 8.f) are those where the electrical artifact visible in figure 4 compromises the low-frequency content of the radar signal. In these maps we can see that for locations further away from $r8$ the correlation with the buoy is much better for the FOBP method than for the SODS method, being able to reach $r25$ with acceptable correlation and RMSE values (see figure 11.b and d). Furthermore, as mentioned above, the interpretation of the RMSE value for the SODS method requires some caution, since it is true that in some locations this statistic is better than for the FOBP method, but this is simply because the SODS method does not capture the variability of the $T_{p}$ signal (Figure 9), being unable to estimate low period values. For this reason, this analysis must be accompanied by the variance difference analysis (Figure 11, e) to understand the RMSE results as reliably as possible.
Figure 11: Same as Figure 8 for $T_{p}$.
4.3 Omnidirectional estimation
A challenging issue is the ability to estimate sea state parameters when the radar time series or Doppler spectra cannot be resolved in azimuth, as is the case for example with compact HF radars. To test this possibility, we applied the SODS and FOBP methods to the different antenna signals without any azimuthal processing. Omnidirectional Doppler spectra and omnidirectional power spectra of amplitude modulation (as defined in section 3.2) were obtained by processing each antenna channel individually and averaging over the $12$ antennas. Due to the variation of the radial flow over the radar coverage, this has the effect of broadening the Doppler spectrum compared to azimuthally resolved Doppler spectra. However, for moderate values of the radial current (which is the case in the Toulon area), this does not cause a critical overlap of the first and second-order regions in the omnidirectional Doppler spectra and the SODS can therefore be applied in the same way. The estimation of $H_{s}$ and $T_{p}$ over the year $2020$ was done in a similar way and the updated statistical power indices are shown in table 2. The evolution of the RMSE of these estimates as a function of the sole range is shown in Figure 12 and the directional estimate at azimuth $40$ and the RMSE from the CANDHIS buoy are given for comparison. It can be seen that the performance of the SODS method is very similar in the directional and omnidirectional cases, with only a slight deterioration in the $H_{s}$ estimate that becomes more pronounced as we move away from the radar location. On the contrary, there is a slight improvement for the $T_{p}$ estimate. This is related to the weak variation of these parameters over the radar coverage, as seen for the WW3 model (Figure 2). The same conclusion applies to the FOBP method for the estimation of $H_{s}$, with a very stable RMSE of about $0.6-0.7$ metres over the range from $8$ to $35$ (i.e. from $12$ to $52.5$ km). Although this RMSE is larger than that obtained with SODS in the short range (about $0.4$ m), it is already better than the latter method in the medium range from $13$ ($19$ km). This is not the case for $T_{p}$, which is very poorly estimated. This could be due to the smearing of the peak frequency of the amplitude modulation $A^{\pm}(t)$ when the radar cell is no longer resolved in azimuth.
Table 2: Same as Table 1 for the omnidirectional SODS and FOBP methods.
$R^{\ast}$ RMSE ($m$ or $s$) $\Delta V$ ($m^{2}$ or $s^{2}$)
SODS FOBP SODS FOBP SODS FOBP
$H_{s}$ r8 $0.90$ $0.73$ $0.44$ $0.60$ $0.058$ $0.085$
r20 $0.82$ $0.64$ $0.83$ $0.63$ $0.408$ $0.117$
r35 $0.42$ $0.71$ $0.010$
$T_{p}$ r8 $0.58$ $0.35$ $1.30$ $1.95$ $1.667$ $1.548$
r20 $0.52$ $0.09$ $1.40$ $2.51$ $1.572$ $2.004$
r35 $0.04$ $2.86$ $2.435$
Figure 12: RMSE of directional (solid lines) and omnidirectional (dashed lines) a) $H_{s}$ in $m$ and b) $T_{p}$ in $s$ from the SODS method (red line) and FOBP method (blue line). The dashed black line corresponds to the mean annual value of $H_{s}$ and $T_{p}$ in the radar coverage. The directional estimation is taken along the azimuth a40.
The scatter-plot of $H_{s}$ estimates from omnidirectional radar measurement is shown in Figure 13 for the same two ranges as considered previously ($r8$ and $r20$) and confirms the robustness of this estimation when no prior selection of bearing can be accomplished, finding the same characteristics as in the directional estimation of each of the methods (see discussion in Section 4.1). It becomes even more evident here that the SODS method overestimates $H_{s}$ as we move further away from the radar.

### v2

@@ 3.2 The Bragg Wave Modulation Method
where $t$ is time; $A^{\pm}$ are complex amplitudes of the form (8) with random phases; $\tilde{f}_{B}$ is the modulated Bragg frequency and $f_{u}=2u/\lambda$ is the additional frequency shift due to the radial surface current $u$. By hypothesis, the variance of the amplitude is proportional to the variance of the long wave steepness:
$\mathrm{\langle\lvert A_{B}^{+}\rvert^{2}\rangle}\sim\mathrm{\langle\lvert k_{p}a_{p}^{\pm}\rvert^{2}\rangle}$ (11)
and thus the same holds for the SWH:
$H_{s}^{2}\sim\mathrm{\langle\lvert A_{B}^{\pm}\rvert^{2}\rangle}$ (12)
The complex Bragg wave amplitudes $A^{\pm}$ are generally different but if we assume that they are uncorrelated due to random phase shifts between the approaching and receding waves, we can simply sum their contribution to the SWH:
$H_{s}^{2}\sim\mathrm{\langle\lvert{A_{B}^{+}}\rvert^{2}\rangle}+\mathrm{\langle\lvert{A_{B}^{-}}\rvert^{2}\rangle}$ (13)
@@
Note that the oscillation factor due to long wave modulation “survives” the radar cell averaging process because the wavelength of the longest waves is not negligible compared to the radar range resolution. For typical scales of $75$ m for the former and $1.5$ km for the latter, about $20$ long waves pass through a given radar cell, so that the mean surface level within that patch oscillates in phase with the long wave with a damping factor of about $1/20$. The amplitude of this damped oscillation remains large compared to the amplitude of the Bragg wavelength (because the ratio of the peak wavelength to the Bragg wavelength is $75/9.3\sim 8$ and the ratio of the corresponding amplitude is $\sim 8^{2}$ for a $k^{-4}$ wave spectrum). In addition, there are also wave groups of comparable or larger size to the radar cell. These groups are responsible for very low frequency modulation of the radar signal. Since both the peak wave frequency $f_{p}$ and the current-induced frequency shift $f_{u}$ are much smaller than the Bragg frequency $f_{B}$, we can rewrite eq. (10) as the sum of an analytical signal $s_{a}^{+}$ (respectively, anti-analytical signal $s_{a}^{-}$) that contains only the positive frequency side (respectively, negative frequency side) of the Bragg peaks:
$s(t)=s_{a}^{+}(t)+s_{a}^{-}(t)$ (14)
@@
$s_{a}^{\pm}(t)\sim A_{B}^{+}(t)e^{\pm 2{\rm{i}}\pi(\tilde{f}_{B}\pm f_{u})t}$ (15)
@@
It follows that the SWH can be extracted from the complex radar signal by taking the absolute value of its analytical and anti-analytical components:
$H_{s}^{2}=C(\mathrm{Var}{(\lvert s_{a}\rvert)}+\mathrm{Var}{(\lvert s_{aa}\rvert)}$ (16)
@@
where $C$ is some calibration factor that must be set for each range and azimuth. In the following, this factor has been calculated for each range cell and azimuth by adjusting the mean value of the right-hand side in (16) $H_{s}$ for the whole $2020$ year to the mean value of the $H_{s}$ obtained with the WWIII model.
According to the heuristic Two-Scale model underlying the modulation approach, the slow variation in amplitude $\lvert A^{\pm}_{B}(t)\rvert$ follows the cycle of the long wave. However, it is clear than in general there is a continuum of scales between the peak wave and other long waves and we expect the slow amplitude to vary according to the peak wave frequency ($f_{p}$). It can be identified at the maximum of the associated power spectra $\sigma_{\lvert s_{a}^{\pm}\rvert}$ (which is defined as the mean square modulus of the Fourier transform of $\lvert s_{a}^{\pm}\rvert$):
$f_{p}=\mathrm{Argmax}(\sigma_{\lvert s_{a}^{+}\rvert}(f)+\sigma_{\lvert s_{a}^{-}\rvert}(f))$ (17)
@@
In practice we calculate the power spectra from $18$ min time series (corresponding to the duration of a WERA SORT file) by applying a Fourier transform on the entire sample and smoothing the spectrum with a $10$ point moving average over frequencies. Figure 7 a) shows an example of the complex radar signal time series recorded on the antenna receiving array after beamforming processing. A slow amplitude modulation is visible, marked by the red lines. Part of this modulation is due to the beat frequency caused by the shift of the surface current from the Bragg frequency (see Guérin and Grilli, (2018)) and another part is due to the long wave groups (eqs. (15)). This last term can be isolated by taking the modulus of the analytical/anti-analytical signal (see figure 7, b)), which has the effect of canceling out the Bragg component and the current-induced frequency shift. The peak wave frequency is sought at the maximum of the corresponding power spectrum (Figure 7, c)) in the range $0.1$ to $0.4$ Hz, which contains the peak wave frequencies for the Mediterranean Sea. Only the positive part of the power spectrum is shown as it is symmetric. Note on Figure 7, b) the very slow modulation at a typical time scale of $100$ sec. This corresponds to very low frequencies (about $0.01$ Hz) that fall far outside of the range of wind-wave gravity waves and could correspond to other phenomena such as infragravity waves.
@@ 4 Results and Discussion
The hourly SWH and peak wave frequency were calculated from the radar data using the above procedure for each radar cell. Due to the azimuthal step of beamforming (1 degree), which is much smaller than the actual angular resolution obtained with a 12 antenna array (1010~{}1010 degrees), the outputs of neighbouring radar cells are highly correlated and we therefore performed a moving average over 5 degrees. Estimations of wave parameters that are outside the permissible range of values expected in the coastal region of Toulon (i.e, such that $H_{s}>5$ m and $f_{p}<0.1$) were eliminated. In addition, an upper threshold was set at $0.35$ Hz for the peak wave frequency calculated with the modulation method, due to the necessary separation of long wave scales. Further outliers were eliminated by removing the nonphysical jumps in the time series (using a threshold of 50 cm and 0.1 sec, respectively, for the permissible change of $H_{s}$ and $f_{p}$ over one hour).
4.1 Statistical performances at the buoy location
The CANDHIS buoy is located at a distance of about 36 km from the radar. Although this relatively large distance is not optimal for estimating the wave parameters from the HFR data, the buoy data can be used to assess the performance of the two proposed methods. Figures 8 and 9 show the time series (re-centered over one month for better visibility) for the SWH and peak wave frequency, respectively, using either the HFR-based estimates, the CANDHIS records or the WWIII model. It appears that both methods give very consistent estimates compared to the buoy measurements, with Barrick’s method appearing to be more accurate but also more lacunary than the Bragg wave modulation method. The plots also suggest that the WWIII estimate is often less accurate than the radar measurements, a fact which is confirmed by a thorough statistical analysis (see further).
To compare the methods in more detail, the radar/buoy scatterplots are shown in Figures 10 and 11 for the annual time series where simultaneous data are available. Along with these plots we calculated the usual statistical parameters, namely the RMSD and the Pearson correlation coefficient (R) (see Table 2). It confirms a better correlation of Barrick’s method with the buoy measurements ($R=0.87$ and $R=0.62$ for the SWH and peak wave frequency, respectively ) while the estimates obtained with the Bragg Wave Modulation Method are more scattered ($R=0.43$ and $R=0.38$, respectively). It also shows that Barrick’s method has a tendency to overestimate both $H_{s}$ and $f_{p}$. This justifies the empirical scaling factor which has been used by several authors (Maresca Jr and Georges,, 1980; Heron and Heron, 1998b, ; Ramos et al.,, 2009) to adjust the radar estimates of $H_{s}$. We found that a scaling factor of about $0.58$ significantly improves the agreement with the buoy and reduces the RMSD from $0.48$ m to $0.29$ m, which is close to the RMSD obtained with the WWIII model ($0.25$ m). Note that this scaling factor is very consistent with the values found in (Maresca Jr and Georges,, 1980) ($0.55$) Ramos et al., (2009) ($0.58$).
Figure 8: Time series of $H_{s}$ (m) from February 1st to April 30, 2020 at the location of the CANDHIS buoy, as estimated by the WWIII model (red lines), Barrick’s method (blue symbols) or the Modulation method (magenta symbols). The buoy measurements are shown with black solid lines.
Figure 9: Same as Figure 8 for the peak wave frequency $fp$ (Hz).
Table 2: Mean bias ($\mu$), RMSD ($\sigma$) and Pierson correlation coefficient ($R$) of Barrick’s method (B), the Modulation method (M) and the WWIII model (W) for the estimation of $H_{s}$ (in meter) and peak wave frequency $f_{p}$ (in Hz) over the year $2020$ by comparison with the CANDHIS buoy measurements (range number $21$ and azimuth number $78$). The RMSD of $H_{s}$ following from Barrick’s estimate is given with ($0.48$ m) and without ($0.29$ m) the scaling factor $0.58$.
$\mu$ $\sigma$ $R$
B M W B M W B M W
$H_{s}$ $0.48$ $-0.11$ $-0.03$ $0.48$ $0.65$ $0.25$ $0.87$ $0.43$ $0.94$
($-0.14$) $(0.29)$
$f_{p}$ $0.01$ $0.006$ $0.004$ $0.043$ $0.059$ $0.037$ $0.62$ $0.38$ $0.70$
@@ B M W B M W B M W
Figure 10: Scatter plot of radar $H_{s}$ and $f_{p}$ from Barrick method versus CANDHIS buoy where each point is colored by the spatial density of nearby points. The thin black line is the regression line with unit slope and the thick dashed line is the regression line with slope $1/0.58$ that shows the need to rescale Barrick’s estimate.